OpenStudy (anonymous):

f(x) = { x2 − 4 _______ x − 2 if x < 2 ax2 − bx + 3 if 2 ≤ x < 3 4x − a + b if x ≥ 3 } What are the values of a and b to make the function continuous?

OpenStudy (anonymous):

way easier than it seems \[\frac{x^2-4}{x-2}=x+2\] if \(x\neq 2\) so the limit at 2 is 4 set \(ax^2-bx+3\) should be 2 when \(x=2\) in other words \[4a-2b+3=2\] or \[4a-2b=-1\]

OpenStudy (anonymous):

=O!

OpenStudy (anonymous):

whoa typo sorry i meant \[4a-2b=+3=4\] so \[4a-2b=1\]

OpenStudy (anonymous):

oh wait i got that part ^

OpenStudy (anonymous):

just didnt know what to do with the bottom 2.

OpenStudy (anonymous):

then replace \(x\) by 3 in the second two equations

OpenStudy (anonymous):

12 - a + b = 9a - 3b + 3

OpenStudy (anonymous):

you should get \[12-a+b=9-3b+3\]

OpenStudy (anonymous):

10a - 4b -9 = 0

OpenStudy (anonymous):

ok right

OpenStudy (anonymous):

or \[9a-4b=9\]

OpenStudy (anonymous):

damn i can't type \[10a-4b=9\]

OpenStudy (anonymous):

so.. do i just set that equal to the other 1..

OpenStudy (anonymous):

\[4a-2b=1\] \[10a-4b=9\] solve for \(a\) and \(b\)

OpenStudy (anonymous):

k thx i got it

OpenStudy (anonymous):

yw